A 26.0 mL sample of a 0.462 M aqueous nitrous acid solution is titrated with a 0.363 M aqueous potassium hydroxide solution. What is the pH at the start of the titration, before any potassium hydroxide has been added?
Answer #1

concentration of HNO2 = 0.462 M

before any addition of KOH :

HNO2 (aq)    ---------------> H+ (aq)   + NO2- (aq)

0.462 M                               0                    0

0.462-x                                x                     x

Ka = x^2 / 0.462 - x

4.5 x 10^-4 = x^2 / 0.462 - x

x = 0.0142

[H+] = 0.0142 M

pH = -log (0.0142)

pH = 1.85